OpenStudy (anonymous):

how can i get x and y intercepts for (f)=4x^2_x+20?

OpenStudy (anonymous):

find out what # is x

OpenStudy (anonymous):

to get the y-intercept, turn the x value and solve for y to get the x-intercept, factor the function, the equal it to 0 and solve for the x

OpenStudy (whpalmer4):

\[f(x) = 4x^2-x+20\] x-intercept is where \(y = f(x) = 0\) y-intercept is where \(y = f(0)\) Clearly, the y-intercept is easy enough to find: \(y = f(0) = 4(0)^2-(0)+20 = \) The x-intercept is a bit trickier. You can factor the equation and set each factor equal to 0, then solve for the value of \(x\); you can use the quadratic formula \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]where the polynomial is in form \(ax^2+bx+c=0\) so here \(a=4, b=-1,c=20\); and there are other methods that probably aren't what your teacher wants you to do. Do you know how to factor \(4x^2-x+20\)?